If the vertical line of a triangle is 24m and the bottom horizontal line is 10m,
then what would the missing length be? Help please & thanks

Answer by checkley77(12844)

IF THE VERTICLE LINE IS ONE OF THE SIDES & IT'S A RIGHT TRIANGLE THEN:
10^2+24^2=C^2
100+576=C^2
C^2=676
C=SQRT676
C=26 m.
iF THE VERTICLE ANGLE IS THE MIDDLE HEIGHT OF THE TRIANGLE THEN:
5^2+24^2=C^2
25+576=C^2
C^2=601
C=SQRT601
C=24.5m.
ALL OTHER VERTICLE LINES WILL RESULT IN THE THIRD SIDE BEING BETWEEN
24.5M. & 26 m.